OpenStudy (anonymous):

Simplify log6 4x + log6 5x + log6 2x

OpenStudy (anonymous):

do you know that there is a theory in log as below?

OpenStudy (anonymous):

\[\log_6 40x^3 \]

sam (.sam.):

\[\huge \log _{6}4x+\log _{6}5x+\log _{6}2x\] \[\huge \log _{6}(4x)(5x)(2x)\] \[\huge \log _{6}(40x ^{3})\]